1.     How do you prepare a 250-mL of 0.125 M NaOH solution? Show your calculations and the methods/steps needed to prepare the solution.

a.      From solid NaOH pellets (by dissolving)

[Calculate the moles needed and convert it to grams]

b.     From a solution of 0.745 M NaOH (by dilution)



1
Expert's answer
2022-01-14T13:48:04-0500

СM = n / V

n (NaOH) = CM x V = 0.125 x 0.25 = 0.031 mol

n = m / M

M (NaOH) = 40 g/mol


a) m (NaOH)solid pellets = n x M = 0.031 x 40 = 1.25 g

b) n1(NaOH) = n2(NaOH)

0.031 = 0.745 x V

V (NaOH) = 0.031 / 0.745 = 0.042 L = 42 ml